OpenStudy (anonymous):

in four hours a hiker in a canyon goes from 892 ft to 256 ft above the canyon floor find the hikers vertival speed

OpenStudy (amistre64):

??

OpenStudy (anonymous):

so how many feet did he go down , if he started at 892 ft and ended up at 256 ft above the canyon floor ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can you calculate it numbers ?

OpenStudy (amistre64):

did he drop at a constant rate? is this the average speed? stuff like that

OpenStudy (amistre64):

\[\frac{892-256}{4}=rate\]

OpenStudy (amistre64):

also known as: \(speed=\cfrac{distance}{thyme}\)

OpenStudy (anonymous):

threw in a bit of spice into the mix ? :)

OpenStudy (amistre64):

lol .... what can i say, its sage advice ;)

OpenStudy (anonymous):

lol .. too bad I can only give one medal,

OpenStudy (anonymous):

lol!! - cries- Please help me with the answer i will study it tomorrow

OpenStudy (anonymous):

(892-256)/4=159

OpenStudy (anonymous):

speed is distance travelled / time taken 892 - 256 =636 vertical speed = 636/4 = 159 ft/ hour

OpenStudy (anonymous):

about thyme you got the answer

OpenStudy (anonymous):

THANK YOU